# The Impact of Train-Test Leakage on Machine Learning-based Android Malware Detection

### v1

@@ Abstract
When machine learning is used for Android malware detection, an app needs to be represented in a numerical format for training and testing. We identify a widespread occurrence of distinct Android apps that have identical or nearly identical app representations. In particular, among app samples in the testing dataset, there can be a significant percentage of apps that have an identical or nearly identical representation to an app in the training dataset. This will lead to a data leakage problem that inflates a machine learning model’s performance as measured on the testing dataset. The data leakage not only could lead to overly optimistic perceptions on the machine learning models’ ability to generalize beyond the data on which they are trained, in some cases it could also lead to qualitatively different conclusions being drawn from the research. We present two case studies to illustrate this impact. In the first case study, the data leakage inflated the performance results but did not impact the overall conclusions made by the researchers in a qualitative way. In the second case study, the data leakage problem would have led to qualitatively different conclusions being drawn from the research. We further propose a leak-aware scheme to construct a machine learning-based Android malware detector, and show that it can improve upon the overall detection performance.
1 Introduction
@@
Machine learning (ML) has been extensively researched for its use in Android malware detection. Generally, in an ML-based vetting system for detecting Android malware, a raw app is first transformed into a suitable numeric representation – a feature vector, a multi-dimensional latent representation, a graph structure, or an image, each tailored to a specific ML model. The transformation is typically a lossy function. As a result, distinct apps may result in identical app representations due to rapid version updates, software component reuse, region-based release of multiple apps, app cloning and repackaging, online no-code app builders that create ready-to-use apps, and so on. In turn, this could lead to a substantial number of app samples in the testing dataset that have representations identical or nearly identical to representations of some samples in the training dataset, a phenomenon known as data leakage [1, 2, 3, 4].
@@
Data leakage has been a recognized phenomenon in ML applications for a long time [5]. Early research defined different types of data leakage [6, 7, 8] which may arise at various stages of ML model development and application. Data leakage has been identified as a significant concern in the implementation of ML algorithms across many scientific fields [1, 2, 3, 4], including in computer security [9]. In this work, we focus on the leakage caused by identical or nearly identical app representations between training and testing datasets, a type of leakage that has been largely unexplored in the security research community. In machine learning, a model learned on a training set shall be tested on data that is disjoint from the training data. If an app in the testing set has a representation that is identical to that of an app in the training, the model is tested on data it has already seen during training, and this is called train-test leakage in machine learning literature. Note that such leakage could arise from the inherent nature of the data rather than errors in experimental design (some call it “inadvertent leakage”). The leakage generally results in inflated performance results that do not reflect the model’s capacity to generalize, since it could simply memorize the leaked data to make a prediction. Some identify the problem as conflating the model’s ability to memorize with its ability to generalize [2]. Even if the representation is not completely identical, but nearly identical, it still calls into question whether the performance measured on the testing data truly reflects the model’s generalizability beyond the data it is trained on. Our framing of the train-test leakage problem follows that used in machine learning literature of domains such as bioinformatics [4], medical imaging [3], and NLP [2].
@@
To answer these questions, we present two case studies of recent research efforts of using ML in Android malware detection. The first is that of Chen et al. [11], which investigated the combination of contrastive and active learning in Android malware detection. The second is from the authors of this paper, where we investigate the application of graph neural networks to Android malware detection. Both works divide training and testing data temporally so that training data pre-date testing data. In both cases, we found that substantial fractions of apps in the testing data have identical or nearly identical representations to at least one app in the corresponding training data. These apps in the testing data are considered leakage based on two different criteria, depending on the specific ML model used.
@@
    For machine learning models that analyze app structure through graph representations, these graphs are first converted into numerical latent representations called “graph embeddings” for the final detection stage. To identify data leakage in these models, we leverage cosine similarity, a well-established metric for comparing vectors. If a testing app’s embedding has a high cosine similarity score with at least one training app’s embedding, we consider the testing app as leakage.
@@
In practice a detector needs to make a prediction for every app, regardless of whether the app is considered leakage or not. Given this fact, does discerning performance measurement differences after removing the leakage app provide any value for informing operations? This is especially important in the second case study – does the new measurement inform operation to use graph neural network in favor of random forest? How does this reconcile with the fact that random forest performs better than graph neural network on the complete testing data – those apps that must be processed any way? To answer this question, we show that by identifying the leakage apps in the testing data, one could build a leak-aware classifier that treats the leaked and non-leaked apps differently. Specifically, if an app is considered a train-test leakage, the leak-aware detector will use the majority vote of the (nearly) identical training apps’ labels to predict the test app’s maliciousness. Otherwise, it uses the machine learned model to predict. Experimental results show that the graph neural network, when used in this leak-aware approach, outperforms the random forest model on the complete testing data. In fact, it outperforms both the stand alone random forest model, and the random forest model used in the same leak-aware approach. Thus, by identifying the leakage data, we reach a qualitatively different conclusion about the advantages of graph neural network compared to that of random forest in practical use situations as well.
In summary, our research investigates the impact of train-test leakage on the evaluation of machine learning-based Android malware detection systems. In general the measured performance drops when the leakage is removed. Identifying and removing such leakage is not only important in showing machine learning’s true capability of generalizing to substantially different Android apps that it has not been trained on, but also has practical implication on how to apply machine learning in this domain. In particular, in identifying the leakage apps, one can design a leak-aware detection scheme that outperforms the machine learned model alone.
The rest of the paper is structured as follows. Section 2 provides some background of ML-based Android malware detection. Section 3 introduces the notion of train-test leakage and evaluation metrics used in this paper. Section 4 and 5 examine two case studies where train-test leakage impacts machine learning based Android malware detection, with different implications for research conclusions. Section 6 introduces the leak-aware detection scheme. Section 7 provides some insights into why so many distinct apps share almost identical representations. We discuss related work in Section 8 and conclude in Section 9.
2 Background
@@
2.1 App Representation for Machine Learning
Vector representation is a common form of input to an ML model. Creating a feature/term vector typically involves extracting specific features from an app, which can be static features, dynamic features, or both. Once the features have been extracted, an app can be transformed into a binary vector, term frequency (TF) vector or a term frequency-inverse document frequency (TF-IDF) vector. Among these, the binary vector is the most widely used in ML-based Android malware detection. Here each vector component corresponds to a feature with $1$ denoting the presence of the feature and 00 denoting its absence. [^1] While one could design feature vectors to use general TF counts/integers or TF-IDF floating numbers, we have not seen much research that adopts these feature representations for Android malware detection. The binary $\{0,1\}$ vectors as described here are used almost universally as representations in the Android malware detection literature.
DL-based Android malware detection has also used vectorized app representations as model inputs [12, 13, 14, 15]. However, DL’s input capability extends beyond vectors, allowing for richer multi-dimensional tensor representations of text, image, graph, and other structured data as inputs. Furthermore, DL has the ability to automatically learn latent vector representations (called embeddings) that capture high-level semantic relationships in its inputs [16, 17]. Prior works have transformed raw bytecode of an APK’s DEX files into a grayscale or color image suitable for convolutional neural networks, which subsequently learn feature representations at many levels [18, 19]. Other works have represented an app as method-level Control Flow Graphs (CFGs) or Data Flow Graphs (DFGs), which can then be fed into convolutional neural networks or graph convolutional network architectures [20, 21, 22].
2.2 Experimental Biases
Despite significant advances in machine learning based Android malware detection, experimental biases and pitfalls persist, demanding careful consideration. Pendlebury et al. [10] highlighted two crucial concerns: 1) spatial bias: test data does not reflect real-world ratio of benign to malicious apps, and 2) temporal bias: apps in training and testing do not follow a realistic timeline. One example of temporal bias is when some apps in the training set are more recent than those in the testing set, essentially using future knowledge to predict the behavior of earlier apps. This bias often occurs when a large amount of data is collected from single or multiple sources and randomly split into training and testing sets without considering the release date of the apps. Models evaluated on such temporally biased data cannot reflect its utility in operation, where such future knowledge is impossible to obtain. Furthermore, Arp et al. [9] identified sampling bias as another pitfall in learning-based security systems, including Android malware detection. This bias occurs when sampling inadvertently introduces artifacts that do not exist in real-world data. For example, if benign apps are primarily collected from one market, while malicious apps are mainly collected from others, the model may end up learning the difference between the two markets instead of the difference between malware and benign apps. Dataset collection needs to be carefully done to avoid these biases.
3 Train-Test Leakage in Android Malware Detection and Metrics for Evaluating its Impact
3.1 Train-Test Leakage
For machine learning models, if the representation of an Android app in the testing set coincides with the representation of another app in the training set, the app in the testing set is considered leakage – the model will be tested on an app that the model was also trained on [^2] Even though the underlying apps are distinct, since an ML model only learns from the app’s representation, identical representations are the same data to the ML model..
Train-test leakage has been extensively discussed for machine learning models applied in other domains, including bioinformatics [4], medical imaging [3], and NLP [2]. It is well known that such data leakage will inflate performance results of a machine learned model, especially if the fraction of leakage is significant. The main problem is that performance metrics measured on the testing set containing leaked data conflates the model’s ability to generalize with its ability to memorize. For the portion in the testing data that are identical to some in training data, the model can simply memorize those data’s label as “prediction” result. This could undermine the validity of conclusions drawn from ML-based research, especially if the conclusion is sensitive to a model’s ability to generalize to data it has not seen in training. In these domains, researchers have addressed the train-test leakage by re-dividing the train/test splits to avoid such leakage. This is not an option for most security problems, and for Android malware detection in particular, since doing so will likely introduce temporal biases as discussed in Section 2.2. To assess the impact of data leakage on evaluation results, we can exclude the “leaked apps” from the testing dataset, re-evaluate the models on the “cleaned” testing data, re-draw all conclusions thereafter, and compare with the results before the leakage removal. To the best of our knowledge, there has not been any study that directly examines train-test leakage’s impact on evaluating machine learning models’ utility in a security problem domain. One recent manuscript [23] identifies train-test leakage in vulnerability datasets for evaluating code language models, but did not isolate the leakage’s impact on evaluating the models’ performance.
Another type of data leakage, which is sometimes confused with train-test leakage, is called temporal leakage [1]. Temporal leakage arises from temporal bias as discussed in Section 2.2. It happens when testing data pre-dates training data, essentially requiring a model to be trained on future data not available at the time of model usage. Temporal leakage is always a flaw in experiment design, since it is evaluating a model under a scenario that can never materialize in real world. Train-test leakage, on the other hand, could very well happen in a realistic model application scenario. For example, due to the nature of Android app evolvement, apps released today may bear similarity to some apps released in the past, creating possible train-test leakage when those older apps are used for training. The problem caused by train-test leakage is the conflation of model’s ability to generalize with its ability to memorize, leading to potentially inaccurate conclusion drawn about a model’s utility in practice. This paper’s sole focus is train-test leakage’s impact. The two case studies we use are both free of temporal leakage.
3.2 Evaluation Metrics
To examine train-test leakage’s impact on machine learning-based model’s performance measurement, we use F1 score, a common metrics adopted by Android malware detetion work, and balanced accuracy, for isolating the impact of train-test leakage on the performance measurement. We briefly explain the rationale below. First some common terminologies:
TP: True Positives (correctly classified malicious apps)
FN: False Negatives (malware mis-classified as benign)
TN: True Negatives (correctly classified benign apps)
FP: False Positives (benign apps mis-classified as malware)
@@ 3.2 Evaluation Metrics
[…]
@@
Precision and F1 score are sensitive to the malware/benign app ratio [10]. When we remove leakage apps from testing set to examine the leakage apps’ impact on the model’s performance measurement, we may also have changed the malware ratio in the testing set, which may impact F1 score by itself. Thus we need metrics that are insensitive to class ratio. One such metric is balanced accuracy (BA), defined as the arithmetic mean of sensitivity and specificity:
$\textit{Sensitivity}=1-\textit{FNR}\hskip 14.45377pt\textit{Specificity}=1-\textit{FPR}$
[…]
Since BA is insensitive to class ratio, we can compare an ML model’s BA on the testing data before and after the leakage is removed. A performance decrease after removing the leakage indicates that the leaked apps in the test data inflated the performance results of the model.
@@
4 Case Study 1: Continuous Learning for Android Malware Detection
@@
4.1 Summary of the Research and Datasets Used
@@
4.2 Identification of Train-Test Leakage
@@
4.3 Impact of the Leakage on Performance Results
@@
Monthly Sample Budget Model Architecture Measurement on Original Data Average Performance (%) Measurement after Leak Removed from Test Average Performance (%)
@@
Monthly Sample Budget Model Architecture Measurement on Original Data Average Performance (%) Measurement after Leak Removed from Test Average Performance (%)
@@
We examined how the presence of train-test leakage might impact the main research conclusion drawn by the authors. We did so by evaluating the models both on the original test data, and on the test data after removing the leakage. We then compared the results. We expected that all models’ performances would drop after leakage in test data is removed. The key question is whether the main conclusion in the paper, that the proposed continuous learning approach outperforms the prior approaches, will still hold after leakage is removed from test data. Before we conducted the experiments, we first ran the model code on the datasets published by the authors to see whether we could replicate similar performance results as reported in Chen et al.’s original paper [11]. Apart from small differences in the performance numbers due to differences in numerical precision across various hardware architectures and non-deterministic operations related to parallelism and threading444 Due to these factors the model might not produce exactly the same results even when run on the same data with the same settings twice., our results replicated those reported in the original paper.
@@
For both APIGraph and AndroZoo datasets, there is a consistent decrease of the overall model performance when leakage is removed from test data. The F1 metrics overall show decrease after leakage is removed from testing data. But F1 metric is also sensitive to the class ratio in the testing data. If the leakage removal reduces the malware ratio in the testing data, this would result in a decreased F1 score even if the model’s predicting capability remains the same. As discussed in Section 3.2, the BA metric isolates the performance results from the change in the malware ratio in the test set. By comparing the BA numbers before and after leakage removal, one can see how much performance is inflated by the train-test leakage. We can see that the two BA metric columns for the same row almost always show a decrease when leak is removed from test. This decrease is more notable for the AndroZoo data (Figure II) than for the APIGraph data (Figure I). This confirms our hypothesis that train-test leakage inflates a machine learning model’s measured performance.
@@
5 Case Study 2: Graph Convolutional Networks for Android Malware Detection

### v2

@@ Abstract
When machine learning is used for Android malware detection, an app needs to be represented in a numerical format for training and testing. We identify a widespread occurrence of distinct Android apps that have identical or nearly identical app representations. In particular, among app samples in the testing dataset, there can be a significant percentage of apps that have an identical or nearly identical representation to an app in the training dataset. This will lead to a data leakage problem that inflates a machine learning model’s performance as measured on the testing dataset. The data leakage not only could lead to overly optimistic perceptions on the machine learning models’ ability to generalize beyond the data on which they are trained, in some cases it could also lead to qualitatively different conclusions being drawn from the research. We present two case studies to illustrate this impact. In the first case study, the data leakage inflated the performance results but did not impact the overall conclusions made by the researchers in a qualitative way. In the second case study, the data leakage problem would have led to qualitatively different conclusions being drawn from the research. We further examine the real-world impact of the data leakage by dissecting the capability of memorization and the capability of generalization of a machine learning model, and show that by removing leakage from testing data, the evaluation results better reflect the machine learning model’s utility in real-world Android malware detection scenarios.
Index Terms:
Android malware detection, Machine learning, Data leakage, Dataset
I Introduction
@@
Machine learning (ML) has been extensively researched for its use in Android malware detection. In such research a raw app is transformed into a suitable numeric representation – a feature vector, a multi-dimensional latent representation, a graph structure, etc., each tailored to a specific ML model. The transformation is typically a lossy function and distinct apps may result in identical app representations. This could lead to a substantial number of app samples in the testing dataset that have representations identical or nearly identical to representations of some samples in the training dataset, a phenomenon known as a type of data leakage [1, 2, 3, 4].
@@
Data leakage has been a recognized phenomenon in ML applications for a long time [5]. Early research defined different types of data leakage [6, 7, 8] which may arise at various stages of ML model development and application. Data leakage has been identified as a significant concern in the implementation of ML algorithms across many scientific fields [1, 9, 2, 3, 4], including in computer security [10]. We focus on the leakage caused by identical or nearly identical app representations between training and testing datasets, a type of leakage that has been largely unexplored in the security research community. In machine learning, a model learned on a training set shall be tested on data that is disjoint from the training data. If an app in the testing set has a representation that is identical to that of an app in the training, the model is tested on data it has already seen during training, and this is called train-test leakage in machine learning literature. Note that such leakage could arise from the inherent nature of the data rather than errors in experimental design. The leakage generally results in inflated performance results that do not reflect the model’s capacity to generalize, since it could simply memorize the leaked data to make a prediction. Some identify the problem as conflating the model’s ability to memorize with its ability to generalize [2]. Even if the representation is not completely identical, but nearly identical, it still calls into question whether the performance measured on the testing data truly reflects the model’s generalizability beyond the data it is trained on. Our framing of the train-test leakage problem follows that used in machine learning literature of domains such as bioinformatics [4], environmental research [9], medical imaging [3], and NLP [2].
@@
To answer these questions, we present two case studies of recent research efforts of using ML in Android malware detection. The first is that of Chen et al. [12], which investigated the combination of contrastive and active learning in Android malware detection. The second is from the authors of this paper, where we investigate the application of graph neural networks to Android malware detection. Both works divide training and testing data temporally so that training data pre-date testing data. In both cases, we found that substantial fractions of apps in the testing data have identical or nearly identical representations to apps in the corresponding training data. These apps in the testing data are considered leakage based on two different criteria, depending on the specific ML model used.
@@
    For machine learning models that analyze app structure through graph representations, these graphs are first converted into numerical latent representations called “graph embeddings” for the final detection stage. To identify data leakage in these models, we leverage cosine similarity, a well-established metric for comparing embeddings. If a testing app’s embedding has a high cosine similarity score with at least one training app’s embedding, we consider the testing app as leakage.
@@
In practice a detector needs to make a prediction for every app, regardless of whether the app has been seen before or not. Given this fact, does discerning performance measurement differences after removing the leakage app provide any beneficial insights for operational decision-making? This is especially important in the second case study – does the new measurement suggest favoring graph neural networks over random forests for operational use? How does this reconcile with the fact that random forests performs better than graph neural network on the complete testing data which comprises all apps necessitating processing?
To answer this question, we observe that in practice, machine learning is just one tool out of many analysts rely on to make a determination on an app’s maliciousness. For apps that are known malware, i.e., already labeled as malware in the training data, there is likely other more efficient approach to detect them, e.g., by using a known signature. Only when a malware is unknown, i.e., corresponding to the non-leakage portion of the testing data, a machine learning model is potentially more useful in providing guidance. Thus testing a machine learning model’s performance on the non-leakage portion of the testing data tells more about the model’s utility in practice.
In summary, our research investigates the impact of train-test leakage on the evaluation of machine learning-based Android malware detection systems. In general the measured performance drops when the leakage is removed. Identifying and removing such leakage is not only important in showing machine learning’s true capability of generalizing to substantially different Android apps that it has not been trained on, but also has practical implication on how to apply machine learning in this domain.
II Background
@@
II-A App Representation for Machine Learning
@@
DL-based Android malware detection has also used vectorized app representations as model inputs [13, 14, 15, 16]. However, DL’s input capability extends beyond one-dimensional, allowing for richer multi-dimensional tensor representations of text, image, graph, and other structured data as inputs. Furthermore, DL has the ability to automatically learn latent vector representations (called embeddings) that capture high-level semantic relationships in its inputs [17, 18]. Prior works have transformed raw bytecode of an APK’s DEX files into a grayscale or color image suitable for convolutional neural networks, which subsequently learn feature representations at many levels [19, 20]. Other works have represented an app as method-level Control Flow Graphs (CFGs) or Data Flow Graphs (DFGs), which can then be fed into convolutional neural networks or graph convolutional network architectures [21, 22, 23].
II-B Experimental Biases
Despite significant advances in machine learning based Android malware detection, experimental biases and pitfalls persist, demanding careful consideration. Pendlebury et al. [11] highlighted two crucial concerns: 1) spatial bias: test data does not reflect real-world ratio of benign to malicious apps, and 2) temporal bias: apps in training and testing do not follow a realistic timeline. One example of temporal bias is when some apps in the training set are more recent than those in the testing set, essentially using future apps to predict the behavior of earlier apps. This bias often occurs when a large amount of data is collected from single or multiple sources and randomly split into training and testing sets without considering the release date of the apps. Models evaluated on such temporally biased data cannot reflect its utility in operation, where such future knowledge is impossible to obtain. Furthermore, Arp et al. [10] identified sampling bias as another pitfall in learning-based security systems, including Android malware detection. This bias occurs when sampling inadvertently introduces artifacts that do not exist in real-world data. For example, if benign apps are primarily collected from one market, while malicious apps are mainly collected from others, the model may end up learning the difference between the two markets instead of the difference between malware and benign apps. Dataset collection needs to be carefully done to avoid these biases.
III Train-Test Leakage in Android Malware Detection
For machine learning models, if the representation of an Android app in the testing set coincides with the representation of another app in the training set, the app in the testing set is considered leakage – the model will be tested on an app that the model was also trained on. Even though the underlying apps are distinct, since an ML model only learns from the app’s representation, identical representations are the same data to the ML model. Train-test leakage has been extensively discussed for machine learning models applied in other domains, including bioinformatics [4], environmental research [9], medical imaging [3], and NLP [2]. It is well known that such data leakage will inflate performance results, especially if the fraction of leakage is significant. The main problem is that performance metrics measured on the testing set containing leaked data conflates the model’s ability to generalize with its ability to memorize. For the portion of apps in the testing data that are identical to some in training data, the model can simply memorize those data’s label as “prediction” result. This could undermine the validity of conclusions drawn from ML-based research, especially if the conclusion is sensitive to a model’s ability to generalize to data it has not seen in training. In these domains, researchers have addressed the train-test leakage by re-dividing the train/test splits to avoid such leakage. This is not an option for most security problems, and for Android malware detection in particular, since doing so will likely introduce temporal biases as discussed in Section II-B. To assess the impact of data leakage on evaluation results, we can exclude the “leaked apps” from the testing dataset, re-evaluate the models on the “cleaned” testing data, re-draw all conclusions thereafter, and compare with the results before the leakage removal. To the best of our knowledge, there has not been any study that directly examines train-test leakage’s impact on evaluating machine learning models’ utility in a security problem domain. One recent manuscript [24] identifies train-test leakage in vulnerability datasets for evaluating code language models, but did not isolate the leakage’s impact on evaluating the models’ performance.
Another type of data leakage, which is sometimes confused with train-test leakage, is called temporal leakage [1]. Temporal leakage arises from temporal bias as discussed in Section II-B. It happens when testing data pre-dates training data, essentially requiring a model to be trained on future data not available at the time of model usage. Temporal leakage is always a flaw in experiment design, since it is evaluating a model under a scenario that can never materialize in real world. Train-test leakage, on the other hand, could very well happen in a realistic model application scenario. For example, due to the nature of Android app evolvement, apps released today may bear similarity to some apps released in the past, creating possible train-test leakage when those older apps are used for training. The problem caused by train-test leakage is the conflation of model’s ability to generalize with its ability to memorize, leading to potentially inaccurate conclusion drawn about a machine learning model’s utility in practice. This paper’s sole focus is train-test leakage. The two case studies we use are both free of temporal leakage.
III-A Evaluation Metrics
@@
F1 score is a common metric adopted by Android malware detetion work, but it is sensitive to the malware/benign app ratio [11]. When we remove leakage apps from testing set to examine the leakage apps’ impact on the model’s performance measurement, we may also have changed the malware ratio in the testing set, which may impact F1 score by itself. For isolating the impact of train-test leakage on the performance measurement, we use balanced accuracy (BA), defined as the arithmetic mean of sensitivity and specificity. The full definitions of the various evaluation metrics can be found in Appendix A. Unlike F1 score, BA is insensitive to class ratio, which will allow us to show the impact of the leakage on the performance measurement. We can compare an ML model’s BA on the testing data before and after the leakage is removed. A performance decrease after removing the leakage indicates that the leaked apps in the test data inflated the performance results of the model.
@@
IV Case Study 1: Continuous Learning for Android Malware Detection
@@
IV-A Summary of the Research and Datasets Used
@@
IV-B Identification of Train-Test Leakage
@@
IV-C Impact of the Leakage on Performance Results
@@
Monthly
Sample
Budget
Model
Architecture
Measurement on Original Data
Average Performance (%)
Measurement after Leak Removed from Test
Average Performance (%)
@@
Monthly
Sample
Budget
Model
Architecture
Measurement on Original Data
Average Performance (%)
Measurement after Leak Removed from Test
Average Performance (%)
@@
We examined how the presence of train-test leakage might impact the main research conclusion drawn by the authors. We did so by evaluating the models both on the original test data, and on the test data after removing the leakage. We then compared the results. We expected that all models’ performances would drop after leakage in test data is removed. The key question is whether the main conclusion in the paper, that the proposed continuous learning approach outperforms the prior approaches, will still hold after leakage is removed from test data. Before we conducted the experiments, we first ran the model code on the datasets published by the authors to see whether we could replicate similar performance results as reported in Chen et al.’s original paper [12]. Apart from small differences in the performance numbers due to differences in numerical precision across various hardware architectures and non-deterministic operations related to parallelism and threading333 Due to these factors the model might not produce exactly the same results even when run on the same data with the same settings twice., our results replicated those reported in the original paper.
@@
For both APIGraph and AndroZoo datasets, there is a consistent decrease of the overall model performance when leakage is removed from test data. The F1 metrics overall show decrease after leakage is removed from testing data. But F1 metric is also sensitive to the class ratio in the testing data. If the leakage removal reduces the malware ratio in the testing data, this would result in a decreased F1 score even if the model’s predicting capability remains the same. As discussed in Section III-A, the BA metric isolates the performance results from the change in the malware ratio in the test set. By comparing the BA numbers before and after leakage removal, one can see how much performance is inflated by the train-test leakage. We can see that the two BA metric columns for the same row almost always show a decrease when leak is removed from test. This decrease is more notable for the AndroZoo data (Table II) than for the APIGraph data (Table I). This confirms our hypothesis that train-test leakage inflates a machine learning model’s measured performance.
@@
V Case Study 2: Graph Convolutional Networks for Android Malware Detection